a boy shouts inside a deep well and hears the echo 0.4s after shouting. if the speed of sound is 340 m/s find the depth the water level in the well
1
Expert's answer
2019-01-21T10:50:29-0500

Let's first find the time that the sound wave needs to travel from the boy to the water level

in the well:

"t = \frac{t_{tot}}{2},"

here,

"t_{tot} = 0.4 s"

is the total time that the sound wave needs to travel from the boy to the

water level in the well and then return back to the boy.

Then, we get:

"t = \frac{t_{tot}}{2} = \frac{0.4 s}{2} = 0.2 s."

Finally, we can find the depth of the water level in the well as follows:

"h = vt = 340 \frac{m}{s} \cdot 0.2 s =68 m."

Answer:

"h = 68 m"

.